OpenStudy (anonymous):

write the equation in slope-intercept form of a line that has a slope of -3 and contains the point (4,-5)

OpenStudy (anonymous):

y = mx + b y = m=slope(x) + b=y intercept. m=slope which is -3 (4,-5) x=4 y=-5 -5 = -3(4) + b Solve from here.